OpenStudy (anonymous):

125 degrees into radians. I know it becomes 2.181 but I need the pi over whatever fraction form. :)

hero (hero):

\\[125^{\circ} \times \frac{\pi}{180^{\circ}}\]

hero (hero):

You want to set it up so that degrees cancel.

OpenStudy (anonymous):

6.8 something?

hero (hero):

Why are you changing it into a decimal. Just reduce the fraction and keep it in fraction form

hero (hero):

Reduce this: \[\frac{125 \pi}{180}\]

OpenStudy (anonymous):

25 / 36 ?

hero (hero):

\[\frac{25 \pi}{36} \space \text{radians}\]

OpenStudy (anonymous):

Thank you!